# Particle Spectra in the Integrated HydroKinetic Model at RHIC BES Energies

### v2

@@ Abstract
We study light-hadron production in Au+Au collisions at $\sqrt{s_{NN}}=7.7$–39 GeV using an extended Integrated HydroKinetic Model (iHKMe). Focusing on transverse momentum spectra, we investigate the sensitivity to key model parameters, in particular the thermalization time scale. We consider two distinct equations of state: one featuring a crossover and the other a first-order phase transition. In both cases, thermalization begins shortly before full nuclear overlap and lasts approximately 1 fm/$c$ across all energies. Both equations of state provide a similarly good description of the soft particle momentum spectra once the other parameters are slightly adjusted. The most pronounced differences arise at the lower RHIC BES energy of $\sqrt{s_{NN}}=7.7$ GeV, particularly in proton and kaon yields, reflecting their sensitivity to the freeze-out parameters.
@@ II.0.1 Initial pre-equilibrium dynamics
The space-time distributions of the macroscopic fields are then constructed as
$T^{\mu\nu}_{\text{urqmd}}(x_{j})$ $=\sum_{i}\frac{p_{i}^{\mu}p_{i}^{\nu}}{p_{i}^{0}}\,{\cal K}_{ij},$ (3)
$J^{\mu}_{\text{urqmd}}(x_{j})$ $=\sum_{i}B_{i}\frac{p_{i}^{\mu}}{p_{i}^{0}}\,{\cal K}_{ij}.$ (4)
Here, $T^{\mu\nu}_{\text{urqmd}}$ represents the stress-energy tensor at the space-time point $x^{\mu}_{j}$, $J^{\mu}_{\text{urqmd}}(x_{j})$ is the baryon current [^3] In this paper, we do not consider separate equations for other conserved charges. The electric charge density is assumed to be proportional to the baryon density, $n_{q}=\frac{Z}{A}n_{b}$, where $\frac{Z}{A}\approx 0.4$ for gold nuclei. The strangeness is always locally zero. , $p^{\mu}_{i}$ denotes the four-momentum of particle $i$, and $B_{i}$ represents its baryon charge. We also utilize a ${\cal K}_{ij}$ for particle smearing to generate a relatively smooth or averaged over the ensemble of similar events tensors.
@@
$\Delta\sigma^{\mu}$ $=n^{\mu}\Delta x\Delta y\Delta\eta,$ (6)
$n^{\mu}$ $=\left(\tau\cosh\eta,0,0,\tau\sinh\eta\right)\,.$ (7)
@@ II.0.2 Thermalzaition
During the thermalization stage, the system is assumed to consist of two coexisting phases. The first is a non-equilibrium component that evolves according to the UrQMD model, as described by Eq. (3). The second corresponds to near-equilibrium matter, which is governed by relativistic hydrodynamics. Its macroscopic tensors are given by
@@
where ${\cal P}_{\tau}={\cal P}(\tau)$ is a weight function such that ${\cal P}(\tau_{0})=1$ at the start of the thermalization stage, ${\cal P}(\tau_{\text{th}})=0$ at the end, and $0<{\cal P}(\tau_{0}<\tau<\tau_{\text{th}})<1$ in between. So this weight function regulates how matter is pumped from one component to another during the thermalization. As in our previous papers [47, 48] we utilize an ansatz for ${\cal P}(\tau)$ inspired by the Boltzmann equation in relaxation time approximation.
@@ III.0.2 Sensitivity of spectra to the parameters
Figure 1: Transverse momentum ($p_{T}$) spectra for $\pi^{-}$, $p$, and $\bar{p}$ (from left to right) averaged over $750$ simulations, with one standard deviation from the mean values, using sets with random parameters within the ranges presented in Table 1. Data correspond to Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV and 20-30% centrality class.
@@ III.0.3 Model finetuning
[…]
@@
The final iHKM parameters for different RHIC BES energies are summarized in Table 2. In Fig. 3, we show how the inferred value of $\tau_{0}$, obtained from pion spectra, depends on the collision energy $\sqrt{s_{NN}}$. We find that $\tau_{0}$ lies within $\pm 0.2$ fm/$c$ of the scaled overlap time, $0.75\,\tau_{\text{overlap}}$. This indicates that, in our model, thermalization begins slightly before the nuclei fully overlap, but still concludes slightly afterward. It is important to note that assuming thermalization starts at a fixed proper time $\tau$ is a simplification, which may be valid only in the central region of the system.
Figure 3: Inferred $\tau_{0}$ values vs nuclei overlapping time from Eq. (1) with $0.75$ factor. The grey shaded zone reflects the estimated uncertainty of $\tau_{0}$ for the crossover equation of state.
Table 2: Optimized iHKM parameters that provide the best description of bulk observables in Au+Au collisions at $\sqrt{s_{NN}}=7.7,11.5,14.5,19.6,27.0$, and $39.0$ GeV. The remaining parameters are fixed at $R=0.5$ fm and $\eta/s=0.08$.
$\sqrt{s_{NN}}$ (GeV) EoS $\tau_{0}$ (fm/$c$) $\tau_{th}$ (fm/$c$) $\epsilon_{\text{sw}}$ (GeV/fm3)
@@ III.0.4 Momentum spectra
Now, we present the iHKM results for transverse momentum spectra in Au+Au collisions at energies of 7.7, 11.5, 14.5, 19.6, 27, and 39 GeV, across two centrality classes: 20-30% and 0-5%, as shown in Figs. 4 and 5. Analysis of the spectra shown in Figs. 5 and 4 reveals several key trends. First, the differences between the results for the two equations of state are relatively small and become negligible at higher collision energies. This trend is also reflected in the model parameters listed in Table 2. The observed behavior can be attributed to the similar description of the QGP phase in both equations of state, while the main differences arise near the transition to the hadron resonance gas. At higher collision energies, the rapid expansion of the system makes this transition occur too quickly for the differences to significantly affect the spectra, whereas at lower energies the differences are more pronounced.
@@
Figure 4: Comparison of iHKM results with STAR data[70, 52] for $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in the 20–30% centrality class at midrapidity ($|y|<0.1$). The pion spectra were used to calibrate the $\tau_{0}$ and $\tau_{\text{th}}$ parameters.
Figure 5: Comparison of iHKM results with STAR data [70, 52] for $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in the 0–5% centrality class at midrapidity ($|y|<0.1$).
@@ IV Summary
Our analysis reveals that the duration of the thermalization stage is approximately 1 fm/$c$ across all RHIC BES energies in the range $\sqrt{s_{NN}}=7.7$$39$ GeV. However, thermalization begins earlier at lower collision energies. Specifically, our numerical results indicate that thermalization commences slightly before the two colliding nuclei fully overlap, assuming they propagate with their initial rapidities; this corresponds to $\tau_{0}\approx 0.75\,\tau_{\text{overlap}}$ for both considered equations of state.
@@ Acknowledgements.
H.Z.’s work was supported by the Grant of National Science Centre, Poland, No: 2021/41/B/ST2/02409 and 2020/38/E/ST2/00019. The research of Yu. S. was partially funded by IDUB-POB projects granted by WUT (Excellence Initiative: Research University (ID-UB)). M.A. gratefully acknowledges support from the Simons Foundation (Grant SFI-PD-Ukraine-00014578).

### v3

# Particle spectra in the integrated hydrokinetic model at RHIC Beam-Energy-Scan energies

@@ Abstract
We study light-hadron production in Au+Au collisions at $\sqrt{s_{NN}}=7.7–39$ GeV using an extended Integrated HydroKinetic Model (iHKMe). Focusing on transverse momentum spectra, we investigate the sensitivity to key model parameters, particularly the thermalization timescale. We consider two distinct equations of state: one featuring a crossover and the other a first-order phase transition. In both cases, thermalization begins shortly before full nuclear overlap and lasts approximately 1 fm/$c$ across all energies. Both equations of state provide a similarly good description of the soft particle momentum spectra once the other parameters are slightly adjusted. The most pronounced differences arise at the lower RHIC BES energy of $\sqrt{s_{NN}}=7.7$ GeV, particularly in proton and kaon yields, reflecting their sensitivity to the freeze-out parameters.
@@ II.0.1 Initial pre-equilibrium dynamics
The space-time distributions of the macroscopic fields are then constructed as {align} T^μν_urqmd (x_j) = ∑_i piμpiνpi0   K_ij,
J^μ_urqmd(x_j) = ∑_i B_i piμpi0   K_ij. Here, $T^{\mu\nu}_{\text{urqmd}}$ represents the stress-energy tensor at the space-time point $x^{\mu}_{j}$, $J^{\mu}_{\text{urqmd}}(x_{j})$ is the baryon current [^3] In this paper, we do not consider separate equations for other conserved charges. The electric charge density is assumed to be proportional to the baryon density, $n_{q}=\frac{Z}{A}n_{b}$, where $\frac{Z}{A}\approx 0.4$ for gold nuclei. The strangeness is always locally zero. , $p^{\mu}_{i}$ denotes the four-momentum of particle $i$, and $B_{i}$ represents its baryon charge. We also utilize a ${\cal K}_{ij}$ for particle smearing to generate a relatively smooth or averaged over the ensemble of similar events tensors.
@@
{align}
Δσ^μ= n^μ Δx Δy Δη,
n^μ = (τcoshη, 0, 0, τsinhη) ,
Here, $\Delta x=\Delta y=0.3$ fm denote the cell size in the transverse plane, and $\Delta\eta=0.05$ is the cell size of the space-time rapidity in the longitudinal direction. We use the same lattice spacing, as well as the proper time step $\Delta\tau=0.05$ fm/$c$, for all energies throughout this paper.
@@ II.0.2 Thermalzaition
During the thermalization stage, the system is assumed to consist of two coexisting phases. The first is a non-equilibrium component that evolves according to the UrQMD model, as described by Eq. (II.0.1). The second corresponds to near-equilibrium matter, which is governed by relativistic hydrodynamics. Its macroscopic tensors are given by
@@
where ${\cal P}_{\tau}={\cal P}(\tau)$ is a weight function such that ${\cal P}(\tau_{0})=1$ at the start of the thermalization stage, ${\cal P}(\tau_{\text{th}})=0$ at the end, and $0<{\cal P}(\tau_{0}<\tau<\tau_{\text{th}})<1$ in between. So this weight function regulates how matter is pumped from one component to another during the thermalization. As in our previous papers [13, 47], we utilize an ansatz for ${\cal P}(\tau)$ inspired by the Boltzmann equation in relaxation time approximation.
@@ III.0.2 Sensitivity of spectra to the parameters
Figure 1: Midrapidity ($|y|<0.1$) transverse momentum spectra $d^{2}N/\left(2\pi p_{T}dp_{T}dy\right)$ for $\pi^{-}$, $p$, and $\bar{p}$ (from left to right), averaged over 750 simulation sets with random parameters within the ranges in Table 1. Shaded regions show one standard deviation from the mean. STAR Collaboration data, used for comparison, correspond to Au+Au collisions at $\sqrt{s_{NN}}=14.5$ GeV in the 20–30% centrality class [8, 4].
@@ III.0.3 Model finetuning
In this study, we calibrate the model’s free parameters using only the particle spectra. As illustrated in Fig. 2, the starting time of the thermalization stage, $\tau_{0}$, is one of the most important parameters. Together with the relaxation time $\tau_{\rm rel}$, they are constrained by the pion multiplicities and the slope of the transverse momentum spectra. In principle, these timescales can depend on the collision centrality. However, our numerical analysis shows that if $\tau_{0}$ is calibrated separately for different centrality classes (between 055% and 30304040%), the optimal values differ by no more than $0.1$ fm/$c$ at $\sqrt{s_{{}_{\rm NN}}}=39$ GeV and by about $0.2$ fm/$c$ at $\sqrt{s_{{}_{\rm NN}}}=7.7$ GeV for both equations of state. Therefore, to reduce the number of free parameters, we present the results obtained for the 20203030% centrality class, where the system is sufficiently large yet exhibits transverse asymmetry due to the nonzero impact parameter.
@@
The final iHKM parameters for different RHIC BES energies are summarized in Table 2. Figure 3 shows the dependence of the extracted $\tau_{0}$ values on the collision energy $\sqrt{s_{NN}}$. We find that $\tau_{0}$ remains close to the nuclei overlap time, $\tau_{\text{overlap}}$, calculated from Eq. (1) and scaled by a factor of 0.75. The deviations do not exceed $\pm 0.2$ fm/$c$, which corresponds to the uncertainty in the calibration of $\tau_{0}$. It should be noted, however, that this overlap time corresponds to zero impact parameter, while our analysis indicates nearly the same $\tau_{0}$ values even for 30304040% non-central collisions, where the nuclei overlap for a shorter duration. This behavior can be attributed to the assumption of uniform thermalization across the system as a function of proper time in our model. In a more realistic scenario, the thermalization rate could depend on the local energy density; however, such an approach would significantly complicate the hydrodynamic equations. Furthermore, the available experimental data for particle spectra are restricted to a narrow midrapidity window, $\left|y\right|<0.1$, which further limits the precision of the parameter extraction.
Another outcome of the model calibration is an estimate of the duration of the thermalization stage. As shown in Table 2, the characteristic scale of $\tau_{\text{rel}}=\tau_{\text{th}}-\tau_{0}$ is of the order of 1 fm/$c$, and no correlation with the collision energy is observed. Likewise, the end of the hydrodynamic stage, characterized by the maximum proper time on the hypersurface of constant energy density $\epsilon_{\text{sw}}$ (hereafter denoted as $\tau_{\text{sw}}^{\max}$ for convenience), exhibits only a weak dependence on the collision energy. For example, in the most central collisions, when decreasing the collision energy from 39 GeV to 7.7 GeV, $\tau_{\text{sw}}^{\max}$ increases only from 9.2 fm/$c$ to 9.8 fm/$c$. For noncentral collisions, this trend is slightly more pronounced. In the 20–30% centrality class, the corresponding times are 8.8 fm/$c$ and 11.5 fm/$c$. Notably, in all these cases, the hydrodynamic stage ends considerably later than the thermalization time $\tau_{\text{th}}$. This indicates that, at the RHIC BES energies considered in this paper, most of the system undergoes thermalization and subsequently evolves through a relatively long hydrodynamic stage. Consequently, the system is largely thermalized, except for the peripheral regions of the fireball with low initial energy density.
Extrapolating these trends to lower collision energies, a qualitative change in the system’s evolution can be anticipated. If $\tau_{0}$ continues to scale with the nuclear overlap time $\tau_{\text{overlap}}$, which increases rapidly in the few-GeV regime, while the end of the hydrodynamic stage $\tau_{\text{sw}}^{\max}$ remains only weakly dependent on energy, full hydrodynamization is not achieved for most of the system. Consequently, particle spectra are expected to contain a substantial non-thermal component. Quantitative estimates indicate that at energies around 4.5 GeV (accessible at STAR FXT [3] and FAIR CBM [10]), where $\tau_{\text{overlap}}\approx 7.5$ fm/$c$, this non-thermal contribution (see Eq. 10) is already significant. At lower energies, near 3 GeV (studied also by HADES [5]), it is expected to dominate. These low-energy regimes will be the focus of our future studies.
Figure 3: Inferred values of the $\tau_{0}$ parameter in iHKM as a function of collision energy $\sqrt{s_{NN}}$ (black markers). The blue line shows the nuclei overlapping time from Eq. (1), scaled by a factor of $0.75$, and the shaded region represents the estimated uncertainty of $\tau_{0}$.
Table 2: Optimized iHKM parameters that provide the best description of transverse momentum spectra in Au+Au collisions at $\sqrt{s_{NN}}=7.7,11.5,14.5,19.6,27.0$, and $39.0$ GeV. The remaining parameters are fixed at $R=0.5$ fm and $\eta/s=0.08$.
$\sqrt{s_{NN}}$ (GeV) EoS $\tau_{0}$ (fm/$c$) $\tau_{\text{th}}$ (fm/$c$) $\epsilon_{\text{sw}}$ (GeV/fm3)
@@ III.0.4 Momentum spectra
We now present the iHKM results for transverse momentum spectra in Au+Au collisions at energies of 7.7, 11.5, 14.5, 19.6, 27, and 39 GeV, for two centrality classes: 20–30% and 0–5%, as shown in Figs. 4 and 5. Analysis of these spectra reveals several key trends. First, the differences between the results obtained with the two equations of state are relatively small and become negligible at higher collision energies. This behavior is also reflected in the model parameters listed in Table 2. The similarity arises from the nearly identical description of the QGP phase in both equations of state, with the main differences occurring near the transition to the hadron resonance gas. At higher collision energies, the rapid expansion of the system makes this transition occur too quickly for the differences to significantly affect the spectra, whereas at lower energies the differences are more pronounced.
@@
Figure 4: Comparison of iHKM results with STAR data [8, 4] for $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ transverse momentum spectra in the 20–30% centrality class at midrapidity ($|y|<0.1$). The pion spectra were used to calibrate the initial time parameters $\tau_{0}$ and $\tau_{\text{th}}$.
Figure 5: Comparison of iHKM results with STAR data [8, 4] for $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ transverse momentum spectra in the 0–5% centrality class at midrapidity ($|y|<0.1$). Model parameters are the same as for the 20–30% centrality class.
@@ IV Summary
Our analysis reveals that the duration of the thermalization stage is approximately 1 fm/$c$ across all RHIC BES energies in the range $\sqrt{s_{NN}}=7.7$$39$ GeV. However, thermalization begins later at lower collision energies. Specifically, our numerical results indicate that thermalization commences slightly before the two colliding nuclei fully overlap, assuming they propagate with their initial rapidities; this corresponds to $\tau_{0}\approx 0.75\,\tau_{\text{overlap}}$ for both considered equations of state.
@@ Acknowledgements.
H.Z.’s work was supported by the Grant of National Science Centre, Poland, No: 2021/41/B/ST2/02409 and 2020/38/E/ST2/00019. The research of Yu. S. was partially funded by IDUB-POB projects granted by WUT (Excellence Initiative: Research University (ID-UB)). The research of M.A. was funded by the National Academy of Sciences of Ukraine through the grant for young scientist research groups. M.A. also gratefully acknowledges support from the Simons Foundation (Grant No. SFI-PD-Ukraine-00014578).